OpenStudy (anonymous):

Each base on a standard baseball diamond lies 90 feet from the next. Find the distance the catcher must throw a baseball from 3 feet behind home plate to second base.

OpenStudy (anonymous):

If the distance between each base is 90 ft. then it makes a square. The distance from home to second is the diagonal of the square. This distance is also the hypotenuse of the triangle formed by home, first, and second. So to find the distance you can either use the pythagorean theorem or just know that you take the side length and multiple by: \[\sqrt{2}\] \[90\sqrt{2}=127.3\] 127.3+3 = 130.3

OpenStudy (anonymous):

Whoa. Thank you. (: